OpenStudy (anonymous):

What is the limit of (x Sin (1/x)) as x becomes positive infinite?

OpenStudy (anonymous):

i believe the limit is 0 because as x increases, 1/x decreases to where it is approaching zero and the sin (0) is 0. and anything multiplied by 0 is 0.

OpenStudy (anonymous):

Use a substitution u=1/x, then the limit will be \[\lim_{u \rightarrow 0}{\sin x \over x}=1\]

OpenStudy (anonymous):

it is 0

OpenStudy (anonymous):

Sorry \[\lim_{u \rightarrow 0} {\sin u \over u}=1\]

OpenStudy (anonymous):

limits is zero , its seen by pinching theorm

OpenStudy (anonymous):

Anwar..i think you might have to use L'Hopital's Rule.

OpenStudy (anonymous):

I don't think L'hopital's rule can work here. I still believe it's 1 by the substitution I used.

OpenStudy (anonymous):

check it, sub in a big number'

OpenStudy (anonymous):

it goes to zero

OpenStudy (anonymous):

the larger you're number gets, the closer you are to zero because it is a fraction.

OpenStudy (anonymous):

@thamir: When substituting \(u=1/x\), \(\sin (1/x)= \sin u\) and \(x=1/u\). Also the approaching point will be zero instead of infinity since \(1/\infty \rightarrow 0\).

OpenStudy (anonymous):

-1<sin(1/x) < 1 , this is known , multiply through by x ( it is positive because we are considering x->+infinity

OpenStudy (anonymous):

-x<xsin(1/x)<x

OpenStudy (anonymous):

wait, thats not going to work :|

OpenStudy (anonymous):

something similar lol

OpenStudy (anonymous):

@elecengineer: When you substitute a large number the x goes to \(\infty \) and the sin(1/x) goes to \(0\). And \( \infty \times 0 \ne0\).

OpenStudy (anonymous):

L'Hopitals Rule does work AnwarA and you are right the answer is 1. lim sin(1/x)/(1/x) =lim [(-1/x^2)cos(1/x)]/(-1/x^2) =lim cos(1/x) =cos(0)=1

myininaya (myininaya):

\[\lim_{x \rightarrow \inf}\frac{\sin(\frac{1}{x})}{\frac{1}{x}}\] let u=1/x x->inf, u->0 so the limit is 1

OpenStudy (anonymous):

Thanks AnwarA for solving it. Thanks chris777 for teaching me the L'Hopitals Rule. And last but not least thanks to elecengineer for trying to solve using the squeeze theorem. Now, I need to get the correct answer which is 1 by using the squeeze theorem. When trying I reached the same point where elecengineer reached. How can we solve it using the squeeze theorem?